# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

### v1

@@ 3 Robust pricing framework
The optimal Bayesian price $p^{*}$ in (2) strongly depends on the (properties of the) hazard rate $f(p)/(1-F(p))$ of some specific valuation distribution. The robust pricing developed in this paper, however, only uses summary statistics and renders a revenue-guarantee accomplished across all possible valuation distributions, without referring to any specific distribution. Therefore, the basic problem for robust monopoly pricing considered in this paper is
@@
with $\mathbb{P}$ some valuation distribution and $\mathcal{P}$ the ambiguity set with all valuation distributions that comply with the partial knowledge that the seller has. Specifically, we consider the ambiguity set (with $\mathcal{B}$ the Borel sigma-algebra over $\mathbb{R_{+}}$)
@@
containing all distributions with mean $\mu$, standard deviation in the interval $[\underaccent{\bar}{\sigma},\bar{\sigma}]$ and support upper bounded by $\beta$. For this ambiguity set, we determine tight bounds for the tail probabilities, solving the minimization part of the central maximin problem (3):
@@
Consider ambiguity set $\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ and denote $\bar{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\bar{\sigma}^{2}}{\beta-\mu}$, $\underaccent{\bar}{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{\beta-\mu}$ and $\underaccent{\bar}{\upsilon}_{2}\mathrel{\mathop{:}}=\mu+\frac{\underaccent{\bar}{\sigma}^{2}}{\mu}$. Then,
@@
Proof (sketch). Lemma 3.1 generalizes the one-sided Chebyshev bound, also known as Cantelli’s inequality, for mean-variance information Feller (1971), and also extends results in De Schepper and Heijnen (1995) for precise knowledge of the variance. Let us sketch the proof of Lemma 3.1. The key step is to formulate the tight bound as the solution to a semi-infinite linear program, see e.g. Shapiro (2001). Writing the tail probability $\mathbb{P}(X>p)$ as the expectation of an indicator function, $\int_{x}\mathds{1}{\{x>p\}}{\rm d}\mathbb{P}(x)$, we see that the tight bound follows from solving the primal problem (with $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$)
@@
by considering all possible shapes of the minorant $F(x)$. For each shape we solve the dual and use complementary slackness to obtain a primal solution, showing that the primal and dual solutions agree. The full proof is presented in Section A. ∎
The solution to (4) in Lemma 3.1 will be leveraged for solving the remaining maximization problem (3) to obtain robust pricing strategies for various choices of information sets that reflect the knowledge of the partially informed seller. In Section 4 we consider the cases where the seller knows either the precise value of the variance or the upper bound on the support (but not both). Building on the results in section 4, we present optimal robust prices in Section 5 for the most general setting considered in this paper, where the seller has the knowledge captured in the ambiguity set (5). In Section 6 we consider a more advanced pricing model for delay-prone services for which we will again employ Lemma 3.1 to perform the maximin analysis.
4 Robust pricing with two prices
This section develops some first robust pricing insights, focusing on two scenarios: when the seller is aware of either the exact value or the maximum value of the standard deviation. We demonstrate that the optimal robust pricing strategy involves choosing one of two prices, as outlined in Propositions 4.1 and 4.2. These propositions play a pivotal role in Section 5, where we extend our analysis to the broader context of knowing the standard deviation falls within a specified interval. In this case, three candidate prices emerge for the optimal pricing strategy.
@@ 4.1 Upper bound on standard deviation
with $\mathcal{P}(\mu,\bar{\sigma},\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and upper bound on standard deviation $\bar{\sigma}$. In this case, the worst-case revenue $\bar{R}(p)$ can thus be viewed as the maximum of $R_{1}(p)$, $R_{2}(p)$ and 00, so the optimal robust price will follow from either $R_{1}(p)$ or $R_{2}(p)$. This is illustrated in Figure LABEL:figmain1. Observe that $\bar{R}(p)$ indeed consists of the two concatenated functions $R_{1}(p)$, $R_{2}(p)$, and that both functions might create a local maximum. For $\bar{\sigma}=0.36$ there are two local maxima, corresponding to two prices, and the global maximum follows from the largest price. As it turns out, the local maximum delivered by $R_{1}(p)$, associated with a relatively low price, is globally optimal when the risk level $\bar{\sigma}$ is sufficiently small. And vice versa, when $\bar{\sigma}$ is sufficiently large, the optimal price is relatively high and corresponds to the argmax of $R_{2}(p)$.
We will now formalize these observations in the first main pricing theorem. To do so, we define the two price functions
$p^{*}_{m}(\beta)$ $=\beta-\sqrt{\beta(\beta-\mu)},$ (10)
and
$p^{*}_{l}(\bar{\sigma})$ $=\mu-\bar{\sigma}\kappa(\mu,\bar{\sigma}),$ (11)
with
The arguments of these price functions are chosen deliberately to express the crucial fact that the price $p^{*}_{l}(\bar{\sigma})$ responds to the risk level $\bar{\sigma}$, while the price $p^{*}_{m}(\beta)$ does not. Here is then our first pricing result, which shows that for a large enough maximal risk-level $\bar{\sigma}$, the optimal price becomes risk-insensitive:
@@
The optimal robust price $\bar{p}^{*}=\arg\sup_{p}\bar{R}(p)$ with $\bar{R}(p)$ in (9) is
@@
with $\bar{\sigma}^{*}=\left(\tfrac{32}{27}(\beta-\mu)\left(\sqrt{\beta(\beta-\mu)}-(\beta-\mu)\right)\right)^{\frac{1}{2}}$.
Proof. The derivative of $R_{1}(p)$ equals zero at $p^{*}_{l}(\bar{\sigma})$ and $\mu$, but only when $p^{*}_{l}(\bar{\sigma})$ lies in the interval $(0,\bar{\upsilon}_{1}]$ is this the maximum ($\mu$ is never the maximum). Since $R_{1}(p)$ increases in $[0,p^{*}_{l}(\bar{\sigma})]$, the maximum is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\min\{p^{*}_{l}(\bar{\sigma}),\bar{\upsilon}_{1}\}$. Similarly, $R_{2}(p)$ is concave and attains a unique maximum at $p^{*}_{m}(\beta)\in[0,\mu]$. Since $R_{2}(p)$ increases in $[0,p^{*}_{m}(\beta)]$, the maximum of $R_{2}(p)$ in $[\bar{\upsilon}_{1},\mu]$ is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\max\{p^{*}_{m}(\beta),\bar{\upsilon}_{1}\}$. We next show that, despite $\bar{\upsilon}_{1}$ being a candidate local maximizer, it can never be the global maximizer for the function $\bar{R}(p)$. Observe that
$R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))$ $=R(p,\mathcal{P}(\mu,\bar{\sigma})\cap\mathcal{P}(\mu,\beta))\geq\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\},$ (14)
and
@@
One important conclusion from Proposition 4.1 is that pricing becomes risk-insensitive when the upper bound on variance is sufficiently large; see Figure LABEL:figmain1. However, this will change drastically when the seller has better or more precise variance information.

### v2

@@ 3 Robust pricing framework
The optimal Bayesian price $p^{*}$ in (1) strongly depends on the properties of the given valuation distribution. The robust pricing developed in this paper, however, only uses summary statistics and renders a revenue-guarantee accomplished across all possible valuation distributions, without referring to any specific distribution. Therefore, the basic problem for robust monopoly pricing considered in this paper is
@@
with $\mathbb{P}$ a valuation distribution from $\mathcal{P}$, the ambiguity set with all valuation distributions that comply with the partial knowledge that the seller has. Specifically, we consider the ambiguity set (with $\mathcal{B}$ the Borel sigma-algebra over $\mathbb{R_{+}}$)
@@
containing all distributions with mean $\mu$, standard deviation in the interval $\underaccent{\bar}{\sigma},\bar{\sigma}]$ and support upper bounded by $\beta$. So instead of knowing an upper bound or exact value, the seller knows that the standard deviation is contained in an interval $\underaccent{\bar}{\sigma},\bar{\sigma}]$. The width of this interval thus reflects how confident the seller is about the actual value $\sigma$, or equivalently, how accurate the seller can estimate $\sigma$.
3.1 Worst-case market
For the ambiguity set (4), we determine tight bounds for the tail probabilities, solving the minimization part (3) of the central maximin problem (2). To this end, we introduce
$\bar{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\bar{\sigma}^{2}}{\beta-\mu},\underaccent{\bar}{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{\beta-\mu},\text{ and }\underaccent{\bar}{\upsilon}_{2}\mathrel{\mathop{:}}=\mu+\frac{\underaccent{\bar}{\sigma}^{2}}{\mu},$ (5)
which serve as threshold values for Lemma 3.1 and play a significant role in the subsequent analysis.
@@ 3.1 Worst-case market
Consider ambiguity set $\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$. Then,
@@
Proof (sketch). Lemma 3.1 generalizes the one-sided Chebyshev bound, also known as Cantelli’s inequality, for mean-variance information (Feller 1971), and also extends results in De Schepper and Heijnen (1995) for precise knowledge of the variance. Let us sketch the proof of Lemma 3.1. The key step is to formulate the tight bound as the solution to a semi-infinite linear program, see e.g. Shapiro (2001). Writing the tail probability $\mathbb{P}(X>p)$ as the expectation of an indicator function, $\int_{x}\mathds{1}{\{x>p\}}{\rm d}\mathbb{P}(x)$, we see that the tight bound follows from solving the primal problem (with $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$)
@@
by considering all possible shapes of the minorant $F(x)$. For each shape we solve the dual and use complementary slackness to obtain a primal solution, showing that the primal and dual solutions agree. The full proof is presented in Section A.1. ∎
From Lemma 3.1 we see that the revenue maximization problem (2) becomes $\sup_{p}\tilde{R}(p)$ with
Let us now examine the structural properties of the worst-case demand function (6) and the corresponding worst-case revenue function (9). The worst-case demand function faced by the robust seller is derived in (6) by selecting, for every fixed price $p$, the valuation distribution that minimizes expected demand. Although this worst-case valuation distribution is a two- or three-point distribution for each fixed $p$, the resulting worst-case demand function is more intricate. To formalize this, we write
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=:1-F^{*}(p),\quad\forall p\in[0,\beta].$ (10)
This representation effectively transforms the distributionally robust setting into the standard Bayesian setting with a known valuation distribution $F^{*}(p)$ defined via the tail bound (6).
Lack of smoothness. One might then consider solving the classic optimization problem $\sup_{p}p(1-F^{*}(p))$ using first-order conditions as in (1). However, this approach fails because the distribution $F^{*}(p)$ does not satisfy the regularity conditions required for a convex revenue function or the existence of a unique optimal price. Figure LABEL:fig:1demrev illustrates $1-F^{*}(p)$ and the resulting revenue function $p(1-F^{*}(p))$ for $\bar{\sigma}=0.3$ across various pairs of $(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$. The left panel reveals that the tail cumulative distribution function (CDF) is non-smooth at certain transition points. At these points, the distribution becomes non-differentiable, leading to non-convexity in the revenue function, as shown in the right panel. Consequently, the minimization problem (3) results in a maximization problem (2) that is non-convex and admits multiple local optima.
Segmented market. Despite the lack of smoothness, the worst-case demand function exhibits a structural shape that aligns intuitively with robust pricing. A relatively large $\bar{\sigma}$ may arise from demand functions associated with heterogeneous customer segments, such as those differing in income, age, or geography. However, the seller possesses only coarse information about market uncertainty in the form of $(\underaccent{\bar}{\sigma},\bar{\sigma})$. If the seller had precise knowledge of distinct market segments and could implement segment-specific pricing, third-degree price discrimination might be an option. Here, we do not consider that possibility; instead, the seller must determine the optimal uniform price. Interestingly, the structure of the worst-case demand function inferred from our analysis suggests a segmented market. The left panel of Figure LABEL:fig:1demrev and Lemma 3.1 indicate that the worst-case market can be interpreted as consisting of three or four segments: consumers willing to pay nothing or prices within intervals $(0,\bar{\upsilon}_{1})$, $(\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1})$, and $(\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2})$. The non-smoothness of $1-F^{*}(p)$ at points like $p=\bar{\upsilon}_{1}$ suggests that as the price surpasses this threshold, a specific segment ceases to purchase the product.
Rise of the niche segment. It is crucial to emphasize that we did not assume a priori the existence of multiple customer groups or a segmented market. Rather, the segmentation emerges as a natural interpretation of the worst-case demand function, with the degree of segmentation increasing as market risk, represented by $\bar{\sigma}$, grows. This phenomenon is evident in Figure LABEL:fig:1demrev, where increasing $\underaccent{\bar}{\sigma}$ amplifies the dominance of the third segment $(\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2})$. Notably, for $\underaccent{\bar}{\sigma}=\bar{\sigma}$, the revenue function exhibits two nearly comparable local maxima. As market risk intensifies, the maximum from the higher segment $(\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2})$ approaches or even surpasses the maximum from the lower segment $(0,\bar{\upsilon}_{1})$, ultimately delivering the global maximum. In Appendix E, we present figures for $\bar{\sigma}$ ranging from 0 to 0.45. These illustrate that for $\bar{\sigma}\geq 0.35$, the higher segment’s local maximum overtakes the lower segment’s maximum. This initial insight—that the robust seller may shift focus from the lower segment to the higher segment in response to increased market risk—will be further substantiated below.
3.2 Main results on robust pricing
We will now present this paper’s main result for robust monopoly pricing. To do so, we introduce three price functions:
$p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ $=\beta-\sqrt{\beta\left(\beta-\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{\mu}\right)},$ (11)
$p^{*}_{m}(\beta)$ $=\beta-\sqrt{\beta(\beta-\mu)},$ (12)
$p^{*}_{l}(\bar{\sigma})$ $=\mu-\bar{\sigma}\kappa(\mu,\bar{\sigma}),$ (13)
with
The arguments of these price functions are chosen deliberately to express the crucial fact that the price $p^{*}_{l}(\bar{\sigma})$ responds to the risk level $\bar{\sigma}$, while the price $p^{*}_{m}(\beta)$ does not. While all three price functions depend on $\mu$, it is always considered fixed and thus omitted as an argument. Notice that unlike $p^{*}_{l}(\bar{\sigma})$ and $p^{*}_{m}(\beta)$ the price function $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ depends on both $\underaccent{\bar}{\sigma}$ and $\beta$. This price function turns out pivotal in high-risk settings.
Theorem 3.1 (Three-Prices Theorem).
with $\Sigma_{l}$, $\Sigma_{m}$ and $\Sigma_{h}$ speficied in (4.3) and such that $\Sigma_{l}\cup\Sigma_{m}\cup\Sigma_{h}$ spanning all pairs $(\underaccent{\bar}{\sigma},\bar{\sigma})$ with $\underaccent{\bar}{\sigma}\leq\bar{\sigma}$.
See Figure 4 for a visual representation of (4.3).
Straight-line boundaries. Figure 4 visualizes Theorem 3.1. The 2D map, divided by straight lines, represents regions corresponding to optimal prices, determined by the market risk bounds—lower and upper—which influence the worst-case demand function (6). For low prices, the demand depends only on the upper bound $\bar{\sigma}$, while for high prices, it is governed by the lower bound $\underaccent{\bar}{\sigma}$. For medium prices, the demand becomes independent of both bounds, making neither significant for determining the optimal price. The boundaries between these regions reflect transitions in how the demand depends on the bounds, whether shifting from one to the other or to neither. This resulting piecewise linear partitioning of the 2D space creates a clear and interpretable map for understanding how market risk bounds influence robust revenue. The oblique line separating the regions corresponding to low and high prices is neither horizontal nor vertical, as each region is determined by a different bound of the standard deviation.
General price jump. When either $\underaccent{\bar}{\sigma}$ or $\bar{\sigma}$ is held constant and does not coincide with the intersection of $\Sigma_{l}$ and $\Sigma_{m}$, varying the other parameter will result in at most two relevant prices, but never three. The high pricing strategy $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ becomes optimal when $\underaccent{\bar}{\sigma}$ and $\bar{\sigma}$ are both sufficiently large. This insight can be described more formally as follows:
Theorem 3.2 (Price-Jump Theorem).
When the upper bound $\bar{\sigma}$ increases from the lower bound $\underaccent{\bar}{\sigma}$ to the maximal value, the optimal price jumps from the low price $p^{*}_{l}(\bar{\sigma})$ to a higher price, which is either $p^{*}_{m}(\beta)$ or $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$, given that $\underaccent{\bar}{\sigma}$ is sufficiently small.
Theorem 3.2 not only identifies the three candidate prices corresponding to the three distinct variance regions but also reveals a critical insight: as demand uncertainty increases, the robust seller strategically introduces discontinuous price jumps to adapt to changing conditions. This insight into the occurrence of discontinuous price adjustments forms the basis of what we term the low-high pricing phenomenon, which serves as the central theme of this paper.
[…]
Figure 4: Three prices theorem for $\mu=0.5$ and $\beta=1$.
The proofs of Theorems 3.1 and 3.2 are provided in Section 4. These proofs build on results for two special cases: knowing the precise value of the variance and knowing the upper bound on the support (but not both). Knowing the true variance corresponds with the diagional in Figure 4, where the lower and upper bound on $\sigma$ agree. The discontinuous price jump for this case can also be seen in Figure LABEL:fig:beta. Knowing only the upper bound corresponds to the vertical axis in Figure 4 with a lower bound of zero. Both cases are analyzed in great detail in Section 4.
3.3 Comparison with ground truth markets
How does truncation influence pricing? When the mean, variance, and support of the distribution are known, we observe that the optimal price follows a non-monotonic pattern—it decreases and then increases as the variance grows. This behavior stands in contrast to the case where only the mean and variance are known. A natural question arises: how does truncation affect pricing behavior? In bounded distributions that result from truncating common unbounded ones—such as truncated exponential, normal, lognormal, or gamma distributions—variance is inherently constrained. While truncation shifts probability mass and alters the shape of the distribution, it does not introduce a fundamental segmentation of the population. Instead, we generally find that the revenue function remains unimodal, and the robust optimal price changes smoothly as variance increases. Figure LABEL:fig:gt_tn shows this monotone behavior for the truncated normal distribution as the ground truth market. The robust price decreases in response to greater dispersion in willingness to pay. In these classical yet truncated smooth markets, we do not observe sharp jumps in pricing; rather, the behavior is primarily driven by variance rather than by an inherent shift in market structure.
Segmentation in beta markets. In contrast, when we consider the beta distribution, a different pattern emerges. Figure LABEL:fig:gt_beta illustrates the performance of the robust price and expected revenue when the actual valuation follows a beta distribution for different values of $\mu$ and $\sigma$. While the scale of expected revenue depends on $\mu$, the relative performance robust pricing is strong and show the universal pattern of low and high pricing. The underlying robust pricing exhibits universal behavior accross all mean values. Unlike truncated distributions, which inherit much of their shape from their unbounded counterparts, the beta distribution can adapt flexibly to any combination of mean, variance, and support. When variance is high, it concentrates most of its probability mass near the minimum and maximum values, forming a distinctive “bathtub” shape. This is a degenerate form of segmentation, where the middle of the distribution is nearly empty, but it does not lead to the bimodal revenue structure seen in the worst-case market. As a result, while the optimal price for the beta distribution follows a pattern similar to that of the worst-case bounded distribution, it never exhibits the discontinuous price jumps observed in the worst-case unbounded setting.
Truncation cannot capture worst-case markets. Ultimately, the key distinction is that the worst-case distribution is inherently bimodal, capturing two distinct groups of consumers—one paying only up to a certain threshold and another willing to pay significantly more. This segmentation drives the sharp transitions in optimal pricing. In contrast, unimodal bounded distributions, whether naturally occurring or arising from truncation, lack this intrinsic segmentation. The beta distribution, though flexible, still results in a unimodal revenue function, ensuring that pricing shifts remain smooth. This fundamental difference in structure explains why the robust optimal price arising from the worst-case market identified in this paper behaves so differently compared to traditional pricing models that assume smooth and regular distributions.
Data-driven comparison. For robust pricing one needs to estimate or learn the key parameters of the ambiguity set. To assess the impact of learning, we assume that the sample of valuation observation grows over time, and we use this growing sample to estimate the mean $\mu$ and maximum valuation $\beta$ via point estimation while constructing a confidence interval for the standard deviation $\sigma$. These four estimates are then used to compute robust prices that results from the matching ambiguity set. We thus consider a setting with $n$ i.i.d. valuations for which the sample mean, sample variance, and maximum valuation
$\bar{X}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i},\quad S_{n}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X}_{n})^{2},\quad M_{n}:=\max\{X_{1},\dots,X_{n}\},$
serve as consistent estimators for $\mu$, $\sigma^{2}$, and $\beta$, respectively. Since $\sigma$ is estimated with uncertainty, we consider as lower and upper bound
[…]
where $\chi^{2}_{\alpha,n-1}$ is the $\alpha$-quantile of the chi-squared distribution with n1n-1 degrees of freedom. Both bounds are consistent estimators of $\sigma$ and are used as lower and upper bounds for $\sigma$ in the ambiguity set, which in turn determines the optimal robust price. Figure LABEL:fig:data-driven illustrates a sample path for the robust price and robust expected revenue when valuations follow a beta distribution with varying $\sigma$. The results indicate that as the standard deviation increases, estimation becomes more challenging but remains feasible, with robust prices and expected revenue still providing meaningful insights. To quantify this effect in more detail, we computed the mean absolute percentage error (MAPE) for both robust price and robust expected revenue. Using 1,000 simulated sample paths, each accumulating data up to $n=20,000$, Table 2 reports MAPE values for several beta distributions. The observed convergence to the true robust quantities supports the reliability of the pricing method in practical applications.
Sample size $n$
Distribution Estimate 100 200 500 1000 2000 5000 10000 20000
($\mu=0.5$, $\sigma=0.2$)
Price 5.73 4.05 2.57 1.85 1.31 0.83 0.57 0.41
$\mathbb{E}(\textup{Revenue})$ 12.36 8.64 5.51 3.97 2.80 1.75 1.23 0.90
Beta / Uniform
($\mu=0.5$, $\sigma=\frac{1}{\sqrt{12}}$)
Price 7.13 5.11 3.30 2.37 1.68 1.06 0.76 0.52
$\mathbb{E}(\textup{Revenue})$ 16.11 11.80 7.52 5.34 3.83 2.44 1.75 1.21
($\mu=0.5$, $\sigma=0.3$)
Price 7.15 5.07 3.37 2.33 1.64 1.01 0.73 0.51
$\mathbb{E}(\textup{Revenue})$ 15.37 11.78 7.94 5.55 3.85 2.41 1.71 1.21
($\mu=0.5$, $\sigma=0.4$)
Price 8.33 5.98 3.91 2.76 2.01 1.23 0.88 0.61
$\mathbb{E}(\textup{Revenue})$ 25.45 19.73 13.79 10.15 7.42 4.81 3.45 2.49
Table 2: Mean absolute percentage error (MAPE) in percentages for various sample sizes and distributions while conducting 1000 simulations.
4 Proofs of robust pricing with two and three prices
This section serves a dual purpose: proving the main pricing theorems and developing further robust pricing insights. We begin by examining two scenarios: one in which the seller knows the exact value (Section 4.1) and another where the seller is aware only of the maximum value of the standard deviation (Section 4.2). In both cases, we show that the optimal robust pricing strategy involves selecting one of two prices, as formalized in Propositions 4.1 and 4.2. These results are instrumental in extending the analysis to the more general setting where the standard deviation lies within a specified interval. The optimal strategy involves three candidate prices, as established in Theorems 3.1 and 3.2, proven in Sections 4.3 and 4.4.
@@ 4.1 Upper bound on standard deviation
with $\bar{\upsilon}_{1}$ as in (5) and $\mathcal{P}(\mu,\bar{\sigma},\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and upper bound on standard deviation $\bar{\sigma}$. In this case, the worst-case revenue $\bar{R}(p)$ can thus be viewed as the maximum of $R_{1}(p)$, $R_{2}(p)$ and 0, so the optimal robust price will follow from either $R_{1}(p)$ or $R_{2}(p)$. This is illustrated in Figure LABEL:figmain1. Observe that $\bar{R}(p)$ indeed consists of the two concatenated functions $R_{1}(p)$, $R_{2}(p)$, and that both functions might create a local maximum. For $\bar{\sigma}=0.36$ there are two local maxima, corresponding to two prices, and the global maximum follows from the largest price. As it turns out, the local maximum delivered by $R_{1}(p)$, associated with a relatively low price, is globally optimal when the risk level $\bar{\sigma}$ is sufficiently small. And vice versa, when $\bar{\sigma}$ is sufficiently large, the optimal price is relatively high and corresponds to the argmax of $R_{2}(p)$.
Here is then a pricing result, saying that for a large enough maximal risk-level $\bar{\sigma}$, the optimal price becomes risk-insensitive:
@@
The optimal robust price $\bar{p}^{*}=\arg\sup_{p}\bar{R}(p)$ with $\bar{R}(p)$ in (16), and $\bar{\sigma}^{*}=\left(\tfrac{32}{27}(\beta-\mu)\left(\sqrt{\beta(\beta-\mu)}-(\beta-\mu)\right)\right)^{\frac{1}{2}}$, is
@@
Proof. The derivative of $R_{1}(p)$ equals zero at $p^{*}_{l}(\bar{\sigma})$ and $\mu$, but only when $p^{*}_{l}(\bar{\sigma})$ lies in the interval $(0,\bar{\upsilon}_{1}]$ is this the maximum ($\mu$ is never the maximum). Since $R_{1}(p)$ increases in $0,p^{*}_{l}(\bar{\sigma})]$, the maximum is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\min\{p^{*}_{l}(\bar{\sigma}),\bar{\upsilon}_{1}\}$. Similarly, $R_{2}(p)$ is concave and attains a unique maximum at $p^{*}_{m}(\beta)\in[0,\mu]$. Since $R_{2}(p)$ increases in $0,p^{*}_{m}(\beta)]$, the maximum of $R_{2}(p)$ in $\bar{\upsilon}_{1},\mu]$ is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\max\{p^{*}_{m}(\beta),\bar{\upsilon}_{1}\}$.
We next show that, despite $\bar{\upsilon}_{1}$ being a candidate local maximizer, it can never be the global maximizer for the function $\bar{R}(p)$. Observe that $R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))=R(p,\mathcal{P}(\mu,\bar{\sigma})\cap\mathcal{P}(\mu,\beta))\geq\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\}$ and
@@
One important conclusion from Proposition 4.1 is that pricing becomes risk-insensitive when the upper bound on standard deviation is sufficiently large; see Figure LABEL:figmain1. However, this will change drastically when the seller has better or more precise standard deviation information.